40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6395     Entry time: Fri Mar 9 16:00:46 2012
Author: steve 
Type: Update 
Category: Green Locking 
Subject: laser emergency shut down switch replaced at the south end 

Over-sized local laser emergency switch was held by large C clamp at the south end. This was replaced by a smaller one and it is mounted with magnets.

The Innolight laser was turned off, while the interlock was wired.

ELOG V3.1.3-